# CaliTex: Geometry-Calibrated Attention for View-Coherent 3D Texture Generation

### v1

@@ Abstract
Figure 1: A collection of 3D objects textured by our method, demonstrating high-fidelity, seamless and geometry-aligned textures facilitated by our framework with geometry-calibrated attention.
@@ 1 Introduction
The automatic generation of 3D assets holds immense promise for content creation in gaming, film, and virtual reality. Recent advances in native 3D generative modeling [39, 50, 21, 53] have significantly improved generation quality, marking a shift from traditional SDS-based pipelines [32, 40] to more powerful diffusion-based approaches. Most advanced techniques adopt a two-stage paradigm: geometry is first synthesized, followed by texture generation conditioned on the geometry. In this process, 2D generative priors from image diffusion models [34, 31, 11, 1, 52] are leveraged to synthesize multi-view images [36, 28, 53], which are then reprojected back onto the surface to construct textures. While this strategy delivers impressive appearance quality, it often collapses under cross-view inconsistency – the same surface region can appear differently across generated views, producing seams or blurring after reprojection. Shown in Fig. 2 (a,c), such artifacts do not stem from rendering, but rather from a deeper representation misalignment within the model itself.
Unlike conventional text-to-image models, where full attention across all noise and text tokens is both sufficient and necessary to capture a semantic context, texture generation involves a far more intricate interplay among noise tokens, geometry condition tokens, and reference image tokens. Current state-of-the-art methods [23, 58, 59] simply extend full attention indiscriminately across all tokens and views, presuming correspondences will emerge. However, this premise does not hold in practice.
@@
Part-Aligned Attention (PAA) mitigates cross-view ambiguity by imposing geometry-aware locality. The key idea is to constrain attention propagation within the same semantic part, preventing tokens from mistakenly attending to geometrically similar but distinct regions across views. Specifically, we use PartField [27] to decompose the mesh into semantic components, assigning each face a discrete part index. By rendering part-colored images from multiple viewpoints, we obtain part labels for every token, enabling us to group tokens that represent the same 3D part across different views. Cross-view attention is then computed within each group, while intra-view full attention is preserved to capture broader context. As shown in Fig. 2 (b), our method effectively eliminates cross-view inconsistency, yielding aligned fine details across different viewpoints.
@@ 2.1 Texture Generation for 3D Objects
With the advent of diffusion-based generative models, texture generation for 3D objects has been extensively explored. Early approaches based on Score Distillation Sampling (SDS) [32, 47, 24, 55, 9, 3, 5] attempted to optimize 3D representations such as NeRF and 3D Gaussian Splatting  [30, 17] by rendering images with sampled random noise and predicting the noise using a pre-trained 2D diffusion model. However, SDS methods often lead to oversaturated colors and suffer from the Janus problem. Recently, the rapid progress of text-conditioned image generation models [11, 20, 25] has inspired many works on 3D texture generation. Some methods [2, 33, 51] iteratively generate single-view images from arbitrary viewpoints, but often suffer from poor cross-view consistency. More recent approaches generate multiple views simultaneously, demonstrating strong quality [6, 15, 21, 38, 23, 59, 50]. However, they fail to achieve proper cross-view and view-geometry alignment when the geometries are highly complex.
Another line of work directly predicts 3D representations with colors [41, 48, 49, 26, 42]. UniTEX [23] introduces a two-stage pipeline: it first generates multi-view images with a diffusion model, followed by a transformer-based Large Texturing Model to predict the texture function in 3D space from both images and geometry. These approaches can inherently avoid issues caused by self-occlusion, but due to the scarcity of large-scale textured 3D training data, their outputs generally underperform compared to methods that leverage existing 2D image generation models in quality.
@@
The cross-view and view-geometry misalignments significantly affect texture quality, and several recent works have attempted to address these issues. SeqTex [50] directly generates UV-space textures alongside multi-view images, replacing the traditional back-projection pipeline. However, recent high-quality 3D generation models [22, 4, 39] typically rely on automatic UV unwrapping tools such as xatlas [46], where fragmented UV layouts often lead to degraded generation quality. Elevate3D [35] refines single-view generations iteratively to improve cross-view consistency and modifies geometry to maintain coherence with synthesized views, though this may inadvertently alter original geometric details. Romantex [12] introduces a 3D-aware RoPE to inject spatial information into the generative model and employs a geometry-related CFG technique to enhance alignment with geometry. AlignTex [56] strengthens geometry alignment by fusing image and geometry features within the diffusion process.
@@
Recent progress in image segmentation, especially with models such as SAM [19], has also driven advances in 3D semantic segmentation. Several methods use 2D segmentation priors from SAM to improve 3D semantic understanding [27, 37, 18, 45, 44]. Among them, PartField [27] trains a network with contrastive learning to predict a continuous 3D feature field, which is then clustered to obtain high-quality 3D segmentation efficiently. In contrast, P3-SAM [29] operates purely in the 3D domain, addressing the spatial inconsistency problem found in methods that depend on 2D segmentation labels. Several works have also explored part-level generation [43, 10] and part-level refinement in 3D generation tasks [4], which inspires us to incorporate semantic part priors into the 3D texture generation.
@@ 3.1 Overview
The overview of our method is shown in Fig. 3. Following previous works [23, 38], we finetune a large-scale image generation network (FLUX.1-Kontext [20]) to synthesize six consistent views from an untextured mesh and a reference image. These generated views are then reprojected to form the texture map and inpainting is applied to fill occluded regions. The backbone of our framework is described in Section 3.2, which consists of two main components: Single-View DiT and Multi-View DiT. In the Single-View DiT stage, we perform batch-wise full attention to capture global information within individual views. In the Multi-View DiT stage, we unfold the batch dimension to perform multi-view attention, enabling information exchange across different views. Our implementations of the modules in Multi-View DiT stage are explained in detail in Section 3.3 and Section 3.4.
@@ 3.2 Multi-View Generation Network
In this section, we describe the general structure of our multi-view generation network based on Flux [1]. Given an untextured mesh and a reference image as inputs, we first render normal maps and canonical coordinate maps (CCMs) from the mesh under six predefined viewpoints. These two maps are averaged and then encoded into the latent space via a variational autoencoder (VAE), producing the geometric condition latent $z_{\mathbf{cond}}$. Similarly, the reference image is encoded into $z_{\mathbf{ref}}$, which is replicated six times to match the batch dimension. Let $z_{t}$ denote the noise latent at timestep $t$. We concatenate the noise, condition, and reference latents along the sequence dimension, forming $\hat{z}_{t}\in\mathbb{R}^{6\times 3L\times C}$, where $L$ and $C$ denote the token length per view and the feature dimension respectively, and the six views are tiled along the batch dimension. Next, a Single-View DiT performs batch-wise full attention to capture intra-view correspondences among the noise, condition, and reference latents.
@@
Finally, the noise latent part $z^{\prime}_{\mathbf{img}}$ is used as the prediction target, and the model is trained via the flow-matching objective $\mathcal{L}(\theta)=\mathbb{E}_{t,z_{0},\epsilon}\!\left[\|z^{\prime}_{\mathbf{img}}-(\epsilon-z_{0})\|^{2}\right]$ [25].
@@ 3.3 Condition-Routed Attention
Next, we describe the two modules in the Multi-View DiT stage. The input to this stage $\tilde{z}_{\mathbf{mv}}\in\mathbb{R}^{1\times 13L\times C}$ consists of three parts: the noise tokens and condition tokens from six views, along with the reference image tokens. Previous approaches [23] simply apply full attention to learn correlations among these modalities. However, this often leads to a cross-modality ambiguity problem, where noise tokens alternately attend to either the reference image or the geometry condition, lacking stable geometric alignment and producing artifacts in the generated textures.
@@ 3.4 Part-Aligned Attention
To address this limitation, we introduce a Part-Aligned Attention (PAA) mechanism that leverages geometric priors to group spatially adjacent tokens across different views. Specifically, we adopt PartField [27] to decompose the 3D mesh $\mathcal{M}$ into $K$ semantic parts (we set $K=20$, which is suitable for most common 3D objects):
@@ 4 Experiments
Step1X-3D [21] 254.1 28.23 3.914 0.8433 0.3154 2.06 2.09 2.07
UniTEX [23] 176.2 17.19 1.156 0.8818 0.3335 3.10 3.21 3.15
MV-Adapter [15] 169.4 13.61 0.747 0.8975 0.2939 3.02 3.03 2.88
Hunyuan3D-2.1 [38] 167.4 16.21 1.067 0.8867 0.3215 2.48 2.60 2.51
@@ 4.1 Implementation Details
We selected 80k objects from Objaverse-XL [7] and Texverse [57] for model training. For each object, we rendered six pre-defined viewpoints at a resolution of 768×768, which were used as the ground truth multi-view images. Additionally, a random view was rendered as the reference image. Our DiT backbone was initialized from FLUX.1-Kontext [20], and we integrate a LoRA adapter [14] with a rank of 16. The model was trained on 8 GPUs for approximately 600 GPU hours in total.
@@ 4.2 Comparisons
We compare our method with several open-source state-of-the-art image-to-texture approaches, including MV-Adapter [15], UniTEX [23], Step1X-3D [21], and Hunyuan3D 2.1 [38]. In addition, we also include comparisons with high-fidelity commercial models, as discussed in Section 4.2.2.
@@
We use FID [13], CLIP-FID, CLIP Maximum-Mean Discrepancy (CMMD) [16], and CLIP-I to evaluate the fidelity of generated textures. We further use LPIPS [54] to measure the perceptual similarity between the generated textures and the ground truth.
@@ 4.2.1 Quantiative Comparison
We evaluate our method on a diverse test set including objects from Objaverse [8] and high-quality game assets, ranging from simple objects to highly detailed models with complex geometry and refined textures. To compare fairly with PBR-based methods, we use illumination-free albedo renderings as reference images. Specifically, for game characters, a random animated pose is sampled as reference while textures are generated on the canonical A-pose to simulate shape discrepancies between reference images and 3D geometries. Each mesh is rendered from 32 viewpoints and compared with the corresponding ground truth albedo images.

### v2

@@ Abstract
Figure 1: A collection of 3D objects textured by our method, demonstrating high-fidelity, seamless and geometry-aligned textures facilitated by our framework with geometry-calibrated attention. Visit our project website at https://calitex-project.github.io.
@@ 1 Introduction
The automatic generation of 3D assets holds immense promise for content creation in gaming, film, and virtual reality. Recent advances in native 3D generative modeling [lai2025hunyuan3d25highfidelity3d, yuan2025seqtexgeneratemeshtextures, li2025step1x, zhang2024clay] have significantly improved generation quality, marking a shift from traditional SDS-based pipelines [poole2022dreamfusion, wang2023prolificdreamer] to more powerful diffusion-based approaches. Most advanced techniques adopt a two-stage paradigm: geometry is first synthesized, followed by texture generation conditioned on the geometry. In this process, 2D generative priors from image diffusion models [rombach2022high, podell2023sdxl, esser2024sd3, blackforest2024flux, zhang2023adding] are leveraged to synthesize multi-view images [shi2023mvdream, long2024wonder3d, zhang2024clay], which are then reprojected back onto the surface to construct textures. While this strategy delivers impressive appearance quality, it often collapses under cross-view inconsistency – the same surface region can appear differently across generated views, producing seams or blurring after reprojection. Shown in Fig. 2 (a,c), such artifacts do not stem from rendering, but rather from a deeper representation misalignment within the model itself.
Unlike conventional text-to-image models, where full attention across all noise and text tokens is both sufficient and necessary to capture a semantic context, texture generation involves a far more intricate interplay among noise tokens, geometry condition tokens, and reference image tokens. Current state-of-the-art methods [liang2025UnitTEX, zhao2025hunyuan3d, zhu2025muma] simply extend full attention indiscriminately across all tokens and views, presuming correspondences will emerge. However, this premise does not hold in practice.
@@
Part-Aligned Attention (PAA) mitigates cross-view ambiguity by imposing geometry-aware locality. The key idea is to constrain attention propagation within the same semantic part, preventing tokens from mistakenly attending to geometrically similar but distinct regions across views. Specifically, we use PartField [liu2025partfieldlearning3dfeature] to decompose the mesh into semantic components, assigning each face a discrete part index. By rendering part-colored images from multiple viewpoints, we obtain part labels for every token, enabling us to group tokens that represent the same 3D part across different views. Cross-view attention is then computed within each group, while intra-view full attention is preserved to capture broader context. As shown in Fig. 2 (b), our method effectively eliminates cross-view inconsistency, yielding aligned fine details across different viewpoints.
@@ 2.1 Texture Generation for 3D Objects
With the advent of diffusion-based generative models, texture generation for 3D objects has been extensively explored. Early approaches based on Score Distillation Sampling (SDS) [poole2022dreamfusion, youwang2024paintit, lin2023magic3d, zhang2024dreammat, deng2024flashtex, chen2023fantasia3d, chen2024textto3dusinggaussiansplatting] attempted to optimize 3D representations such as NeRF and 3D Gaussian Splatting  [mildenhall2020nerfrepresentingscenesneural, kerbl20233dgaussiansplattingrealtime] by rendering images with sampled random noise and predicting the noise using a pre-trained 2D diffusion model. However, SDS methods often lead to oversaturated colors and suffer from the Janus problem. Recently, the rapid progress of text-conditioned image generation models [esser2024sd3, labs2025flux1kontextflowmatching, lipman2023flowmatchinggenerativemodeling] has inspired many works on 3D texture generation. Some methods [chen2023text2tex, richardson2023texturetextguidedtexturing3d, zeng2024paint3d] iteratively generate single-view images from arbitrary viewpoints, but often suffer from poor cross-view consistency. More recent approaches generate multiple views simultaneously, demonstrating strong quality [cheng2024mvpaint, huang2025mv, li2025step1x, hunyuan3d2025hunyuan3d2.1, liang2025UnitTEX, zhu2025muma, yuan2025seqtexgeneratemeshtextures, bao2025lumitex]. However, they fail to achieve proper cross-view and view-geometry alignment when the geometries are highly complex.
Another line of work directly predicts 3D representations with colors [xiang2025structured3dlatentsscalable, yu2023texturegeneration3dmeshes, yu2024texgen, 10656692, xiong2025texgaussiangeneratinghighqualitypbr]. UniTEX [liang2025UnitTEX] introduces a two-stage pipeline: it first generates multi-view images with a diffusion model, followed by a transformer-based Large Texturing Model to predict the texture function in 3D space from both images and geometry. These approaches can inherently avoid issues caused by self-occlusion, but due to the scarcity of large-scale textured 3D training data, their outputs generally underperform compared to methods that leverage existing 2D image generation models in quality.
@@
The cross-view and view-geometry misalignments significantly affect texture quality, and several recent works have attempted to address these issues. SeqTex [yuan2025seqtexgeneratemeshtextures] directly generates UV-space textures alongside multi-view images, replacing the traditional back-projection pipeline. However, recent high-quality 3D generation models [li2025sparc3dsparserepresentationconstruction, chen2025ultra3defficienthighfidelity3d, lai2025hunyuan3d25highfidelity3d] typically rely on automatic UV unwrapping tools such as xatlas [xatlas], where fragmented UV layouts often lead to degraded generation quality. Elevate3D [Ryu_2025] refines single-view generations iteratively to improve cross-view consistency and modifies geometry to maintain coherence with synthesized views, though this may inadvertently alter original geometric details. Romantex [feng2025romantex] introduces a 3D-aware RoPE to inject spatial information into the generative model and employs a geometry-related CFG technique to enhance alignment with geometry. AlignTex [zhang2025aligntex] strengthens geometry alignment by fusing image and geometry features within the diffusion process.
@@
Recent progress in image segmentation, especially with models such as SAM [Kirillov_2023_ICCV], has also driven advances in 3D semantic segmentation. Several methods use 2D segmentation priors from SAM to improve 3D semantic understanding [liu2025partfieldlearning3dfeature, tang2025segmentmesh, kim2024garfieldgroupradiancefields, ying2023omniseg3domniversal3dsegmentation, ye2024gaussiangroupingsegmentedit]. Among them, PartField [liu2025partfieldlearning3dfeature] trains a network with contrastive learning to predict a continuous 3D feature field, which is then clustered to obtain high-quality 3D segmentation efficiently. In contrast, P3-SAM [ma2025p3samnative3dsegmentation] operates purely in the 3D domain, addressing the spatial inconsistency problem found in methods that depend on 2D segmentation labels. Several works have also explored part-level generation [yang2025omnipartpartaware3dgeneration, ding2025fullpartgenerating3dresolution] and part-level refinement in 3D generation tasks [chen2025ultra3defficienthighfidelity3d], which inspires us to incorporate semantic part priors into the 3D texture generation.
@@ 3.1 Overview
The overview of our method is shown in Fig. 3. Following previous works [liang2025UnitTEX, hunyuan3d2025hunyuan3d2.1], we finetune a large-scale image generation network (FLUX.1-Kontext [labs2025flux1kontextflowmatching]) to synthesize six consistent views from an untextured mesh and a reference image. These generated views are then reprojected to form the texture map and inpainting is applied to fill occluded regions. The backbone of our framework is described in Section 3.2, which consists of two main components: Single-View DiT and Multi-View DiT. In the Single-View DiT stage, we perform batch-wise full attention to capture global information within individual views. In the Multi-View DiT stage, we unfold the batch dimension to perform multi-view attention, enabling information exchange across different views. Our implementations of the modules in Multi-View DiT stage are explained in detail in Section 3.3 and Section 3.4.
@@ 3.2 Multi-View Generation Network
In this section, we describe the general structure of our multi-view generation network based on Flux [blackforest2024flux]. Given an untextured mesh and a reference image as inputs, we first render normal maps and canonical coordinate maps (CCMs) from the mesh under six predefined viewpoints. These two maps are averaged and then encoded into the latent space via a variational autoencoder (VAE), producing the geometric condition latent $z_{\mathbf{cond}}$. Similarly, the reference image is encoded into $z_{\mathbf{ref}}$, which is replicated six times to match the batch dimension. Let $z_{t}$ denote the noise latent at timestep $t$. We concatenate the noise, condition, and reference latents along the sequence dimension, forming $\hat{z}_{t}\in\mathbb{R}^{6\times 3L\times C}$, where $L$ and $C$ denote the token length per view and the feature dimension respectively, and the six views are tiled along the batch dimension. Next, a Single-View DiT performs batch-wise full attention to capture intra-view correspondences among the noise, condition, and reference latents.
@@
Finally, the noise latent part $z^{\prime}_{\mathbf{img}}$ is used as the prediction target, and the model is trained via the flow-matching objective $\mathcal{L}(\theta)=\mathbb{E}_{t,z_{0},\epsilon}\!\left[\|z^{\prime}_{\mathbf{img}}-(\epsilon-z_{0})\|^{2}\right]$ [lipman2023flowmatchinggenerativemodeling].
@@ 3.3 Condition-Routed Attention
Next, we describe the two modules in the Multi-View DiT stage. The input to this stage $\tilde{z}_{\mathbf{mv}}\in\mathbb{R}^{1\times 13L\times C}$ consists of three parts: the noise tokens and condition tokens from six views, along with the reference image tokens. Previous approaches [liang2025UnitTEX, bao2025lumitex] simply apply full attention to learn correlations among these modalities. However, this often leads to a cross-modality ambiguity problem, where noise tokens alternately attend to either the reference image or the geometry condition, lacking stable geometric alignment and producing artifacts in the generated textures.
@@ 3.4 Part-Aligned Attention
To address this limitation, we introduce a Part-Aligned Attention (PAA) mechanism that leverages geometric priors to group spatially adjacent tokens across different views. Specifically, we adopt PartField [liu2025partfieldlearning3dfeature] to decompose the 3D mesh $\mathcal{M}$ into $K$ semantic parts (we set $K=20$, which is suitable for most common 3D objects):
@@ 4 Experiments
Step1X-3D [li2025step1x] 254.1 28.23 3.914 0.8433 0.3154 2.06 2.09 2.07
UniTEX [liang2025UnitTEX] 176.2 17.19 1.156 0.8818 0.3335 3.10 3.21 3.15
MV-Adapter [huang2025mv] 169.4 13.61 0.747 0.8975 0.2939 3.02 3.03 2.88
Hunyuan3D-2.1 [hunyuan3d2025hunyuan3d2.1] 167.4 16.21 1.067 0.8867 0.3215 2.48 2.60 2.51
@@ 4.1 Implementation Details
We selected 80k objects from Objaverse-XL [deitke2023objaversexl] and Texverse [zhang2025texverseuniverse3dobjects] for model training. For each object, we rendered six pre-defined viewpoints at a resolution of 768×768, which were used as the ground truth multi-view images. Additionally, a random view was rendered as the reference image. The back-projection and inpainting modules are adopted from Lumitex [bao2025lumitex]. Our DiT backbone was initialized from FLUX.1-Kontext [labs2025flux1kontextflowmatching], and we integrate a LoRA adapter [hu2022lora] with a rank of 16. The model was trained on 8 GPUs for approximately 600 GPU hours in total.
@@ 4.2 Comparisons
We compare our method with several open-source state-of-the-art image-to-texture approaches, including MV-Adapter [huang2025mv], UniTEX [liang2025UnitTEX], Step1X-3D [li2025step1x], and Hunyuan3D 2.1 [hunyuan3d2025hunyuan3d2.1]. In addition, we also include comparisons with high-fidelity commercial models, as discussed in Section 4.2.2.
@@
We use FID [heusel2017fid], CLIP-FID, CLIP Maximum-Mean Discrepancy (CMMD) [jayasumana2024cmmd], and CLIP-I to evaluate the fidelity of generated textures. We further use LPIPS [zhang2018lpips] to measure the perceptual similarity between the generated textures and the ground truth.
@@ 4.2.1 Quantiative Comparison
We evaluate our method on a diverse test set including objects from Objaverse [deitke2023objaverse] and high-quality game assets, ranging from simple objects to highly detailed models with complex geometry and refined textures. To compare fairly with PBR-based methods, we use illumination-free albedo renderings as reference images. Specifically, for game characters, a random animated pose is sampled as reference while textures are generated on the canonical A-pose to simulate shape discrepancies between reference images and 3D geometries. Each mesh is rendered from 32 viewpoints and compared with the corresponding ground truth albedo images.